In a gas mixture, which component will have the higher partial pressure? the one with the higher mole number or the one with the larger molecular weight?
Answer #1

PLEASE .. Please appreciate the work doneby UPVOTING ... Thank you .

In any Gas Mixture ,

Partial Pressure = Mole fraction x Total Pressure .

Now , mole fraction = n/nt ; and n = m / M .

Thus Partial pressure will be higher for higher mole and will be lowered for higher molecular weight , since M = Molecular weight .

Thus Partial pressure will be higher when mole no will be higher